Three of the adult participants ran the top three farthest distances. The first ran 9 1/2km, the second ran 8 1/5km, and the last one ran 8 1/7km. Find the total number of kilometers they ran altogether.
Answer by checkley77(12844)

9 1/2+8 1/5+8 1/7
19/2+41/5+57/7 (LCD=70)
(19*35+41*14+57*10)/70
(665+574+570)/70
1,809/70=25 59/70 ans.